OpenStudy (anonymous):

is e^(xy) +y =x-1 an implicit solution to dy/dx = (e^(-xy) -y))/(e^(-xy) +x)) ?

OpenStudy (anonymous):

use implicit diff. on e^(xy) +y =x-1 see if it matches the RHS

OpenStudy (anonymous):

Or I should say.... use implicit diff. on e^(xy) +y =x-1 sub.s in for dy/dx into the other expression...

OpenStudy (anonymous):

that si swhat i tried but i am getting a wiered answer

OpenStudy (anonymous):

i have dy/dx =-ye^(xy) / x e^(xy)

OpenStudy (anonymous):

let me check...

OpenStudy (anonymous):

hmm I got a different result

OpenStudy (anonymous):

what do you get

OpenStudy (anonymous):

ok I got the correct result... I just checked..

OpenStudy (anonymous):

look closer at your differentiation of \[e ^{xy}\]

OpenStudy (anonymous):

remember that \[(e ^{f(x)})\prime = e ^{f(x)} * f \prime (x)\]

OpenStudy (anonymous):

that's a bit messy... let me use the draw pad

OpenStudy (anonymous):

ok ...alright

OpenStudy (anonymous):